Astronomy Report – January 16th

Astronomy Report

Name: Robinson Raphael

Crew: 219

Date: Jan 16th, 2020

MDRS ROBOTIC OBSERVATORY: (MDRS-14)

Submitted an observation for the Owl Nebula (aiming for longer exposure time)

Images submitted with the report: (See attached below)

Whirlpool Galaxy (Generic and Johnson-Cousins Filters Used)

M40 (Generic and Johnson-Cousins Filters Used)

NGC 4258 (Generic and Johnson-Cousins Filters Used)

Problems encountered: Using the generic filters tend to leave the final image towards one more dominant color.

Science Report – January 16th

Crew 219 Science Report 16-JAN-20

Crew Science Officer: Hannah Blackburn

1. Biometrics and Neurobehavioral Research

All sleep logs and post-EVA surveys sent to HSO.

2. UAV

Disassembled drone for transport and discovered that multiple wires were not connected to motors 3 and 4.

3. Dust Mitigation for Optical Mirrors

Mirrors packed for transport.

4. Astrophotography of Celestial Bodies

Made color images of NGC4258, M40, and the whirlpool galaxy.

5. Remediation of Mars Regolith

Nothing to report.

6. Chemical and Mineralogical Composition of the MDRS Site

Samples prepared for transport.

7. Protocols for the Discovery of Life on Mars

Samples photographed and prepared for transport.

Glassware check out:

10 glass beakers (sizes between 200 and 500 ml) were cleaned, dried, and put away.

Greenhab Report – January16th

Crew 219 GreenHab Report 16-JAN-20

Crew GreenHab Officer: Nathan Hadland

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 23.7 °C

8:09 A.M.
Floor Unit: 18 °C
Mid-height unit: 23 °C

10:31 A.M.
Floor Unit: 19 °C
Mid-height unit: 23 °C

12:24 P.M.
Floor Unit: 20 °C
Mid-height unit: 30 °C

2:41 P.M.
Floor Unit: 12 °C
Mid-height unit: 21 °C

5:08 P.M.
Floor Unit: 16 °C
Mid-height unit: 21 °C

Hours of supplemental light: Light system 7:00 PM-12:00 AM.

Daily water usage for crops: 3 Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – 149.03 Gal.

Time(s) of watering for crops:

10:31 A.M.

SEASONAL PLANTING HISTORY:

Change to crops: N/A

OTHER NOTES:

Harvest: At 10:31 A.M., 2 g of basil, 2 g of oregano, 1 g of thyme, and 1 g of marjoram were harvested

Other: N/A

Journalist Report – January 16th

CREW 219 JOURNALIST REPORT 16JAN2020

AUTHOR: Keith Crisman

SOL 11

ALL THE FEELS FROM MARS: ALPHA AND OMEGA AND ALPHA AGAIN

The beginning of the end… and rebirth. Last night, with little

the ceremony, roles among the crew shifted. As Crew 219’s mission begins

nearing its termination point they will begin to prepare the Hab for

the arrival of their replacement crew and transport back to Earth.

Further, training for ARES’ next season MDRS crew’s CO (C. Montanez,

current GHO) and XO (K. Crisman, current HSO) begins with a transfer

of titles from the current CO (D. Masaitis) and XO (N. Hadland). Our

new (acting) CO and XO are taking the reins with planning and

execution of duties and responsibilities of the roles they will

proudly inhabit next season. As such, in a combination of the end and

a new beginning; today marked the last day our crew held EVAs, the

very last of which was commanded by our (acting) CO and XO.

These EVAs were a slower pace, returning superfluous samples, giving the

crew time to slow down and reflect on the environment, landscape, and

the natural beauty of our Martian desert home. For some, they are ready

to go home, full of stories and experiences ready to share with their

family, friends, and peers. For others such as myself, despite the

short two-weeks stay, and badly missing my wife, daughter, and home

(not to mention a good steak, hot shower, and soft bed), I can’t help

but look out over the sheer beauty of this landscape. It may have

been a short stay, but this crew has become family, this desert has

become home, and although I am ready to leave, part of me will forever

stay here.

To my crew mates and those that follow our stories; Per Scientiam ad Martis, Semper Exploro! To Mars through Science, Always Exploring!

Operations Report – January 16th

Crew 219 Operations Report 16-JAN-2020

Sol: 11
Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run
Hours run: 14
From what time last night: 17:30
To what time this morning: 07:30
List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 80%

09:10 = 100%; 10:00 = 99%;
11:00 = 97%;
13:00 = 88%;
14:00 = 85%;
15:00 = 82%;
16:30 = 80%;
18:00 = 100%

Diesel Reading – 55%
Station Propane Reading- 32%
Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50
Water Meter (units)- 0147767,3
Water (static tank) (gal) – 286
Static to loft Pump used –Yes
Water in Green Hab (gal): 149.03
Water in Science Dome (gal): 0

Toilet Tank Emptied –Yes

Deimos rover used: Still in the Workshop

Hours:
Beginning Charge:
End Charge:
Currently Charging:

Sojourner rover used: Assigned to Director
Hours:
Beginning Charge:
End Charge:
Currently Charging:

Spirit Rover used- Yes
Starting Hours: 126.1, 126.2
Beginning Charge: 100%, 100%
Ending Hours: 126.2, 126.4
Ending Charge: 51%, 76%

Opportunity Rover used: Still in the workshop
Hours:
Beginning Charge:
Ending Charge:
Currently Charging:

Curiosity Rover used: Yes
Starting Hours: 132.0, 132.1
Beginning Charge: 100%, 100%
Ending Hours: 132.1, 132.4
Ending Charge: 51%, 78%

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use:
Oil Added? No
# hours the ATV’s were used today:
Notes on ATVs:

Habcar used and why, where?
CrewCar used and why, where? In the shop.

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: Suit #7 began a rapid power drop halfway through EVA 17.

Summary of Hab: Nominal

Summary of Science Dome operations: Average daytime temperature is 7°C. The average night time temperature is 2°C.

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support:

EVA Report – January 16th

Crew 219 EVA Report 16-01-2020

EVA #17

Author: Dave Masaitis

Purpose of EVA: Sample return

Start time: 0908 hrs
End time: 1013 hrs

Narrative: EVA #17 completed a five-minute decompression by 0908 hours, mounted Spirit and Curiosity, and departed the Hab campus after reporting hours and SOC. The rovers proceeded northbound on Cow Dung Road until achieving a positive ID of Streambed Connector. The EVA team parked rovers and dismounted, proceeding east on foot along Streambed Connector. The streambed was full of patchy snow and ice, so the team moved cautiously along their directed route.

The route added approximately thirty minutes to planned transit time, so the team had to move swiftly to deposit samples once at their destination. Once complete, they discussed their route back and opted to try and retrace their incoming route as closely as possible. Only one team member fell through the ice, but no injuries were sustained, so they continued their movement to the rovers. The team proceeded southbound on Cow Dung Road to the Hab, re-entering the airlock at 1013 hours.

Destination: West of Candor Chasma
Coordinates: 0520000/4251500

Participants: D. Masaitis (HSO), N. Hadland (GHO), A. Elnajdi (GEO)

Roads and routes per MDRS Map: North on Cow Dung Road to Streambed Connector, East through Streambed Connector

Mode of travel: Rovers (Spirit and Curiosity) and Walking

EVA #18

Author: Cynthia Montanez

Purpose of EVA: Sample Return

Start Time: 1407 hrs
End Time: 1539 hrs

Narrative: EVA #18 is the last EVA crew 219 was able to do before completing their Mars analog mission. At 1407 hours, Cynthia (acting as CO) and Keith (acting as XO) lead their team 1 km south of Kissing Camel on Curiosity as Dave and Alejandro followed on Spirit.

When reaching their destination, Dave directed his crew to the specific locations each sample was taken from so that they could return their samples. Once completed, the EVA crew made their way back to the rovers and proceeded safely back to the Hab at 1539 hours.

Destination: 1 km S of Kissing Camel
Coordinates: 0518500/4248300

Participants: Cynthia Montanez (CO), Keith Crisman (XO), Dave Masaitis (HSO), Alejandro Perez (ENG)

Roads and Routes per MDRS Map: South on Cow Dung Road (0110), approximately 1 km south of Kissing Camel to return samples.

Mode of travel: Rovers (Spirit and Curiosity) and walking.

Sol Summary – January 16th

Crew 219 Sol Summary Report 16-01-2020

Sol: 11
Summary Title: If You Give a Martian a Cookie
Author’s name: Cynthia Montanez

Mission Status: Nominal

Sol Activity Summary: The morning started with crew 219 sipping on coffee and indulging in a light breakfast before they partook in their nominal tasks. To train C. Montanez and K. Crisman for their next field season, C. Montanez took the position as CO and K. Crisman took the position as XO for the day. Between 0908 to 1018, EVA #17 entered sim for sample return and came back to a delicious dish of pork sausage meatballs and spaghetti made by H. Blackburn.

Once everyone finished eating, each person started to work on finishing up their projects, post-mission report edits, and/or cleaning up parts of the Hab. Between 1402 to 1534, EVA #18 entered sim for the final sample return and came back to freshly made chocolate chip oatmeal cookies, brownies, and chai tea made by H. Blackburn. The aroma of the desserts made the Hab smell like grandma’s house, as the taste brought the crew back to Earth.

Finally, the day ended with H. Blackburn cooking coconut curry and crew positions being switched back. Overall, today was a great day on Mars.

Look Ahead Plan: Cleaning the Hab and its exterior facilities in preparation for the next crew

Weather: Chilly and sunny all-day

Crew Physical Status: Nominal

EVA:

EVA #17 – Sample Return W of Candor Chasma (0520000/4251500)
EVA #18 – Sample Return 1 km S of Kissing Camel (0518500/4248300)

Reports to be filed:

Sol Summary
EVA Reports #17 and #18
Operations Report
GreenHab Report
Research Report
Astronomy Report
Journalist’s Report
Photos (Including Photo of the Day)

Support Requested: None